4) elect Shot 2019-01-28 at 8.36.40 PM htt 99 S ve Question 8 (1 point) Two equally charged spheres of mass 1.0 g are placed 2.0 cm apart. When released, they begin to accelerate at 414 m/s2. What is the magnitude of the charge on each sphere? OA) 75 nG B) 140 nC c) 95 nC D) 120 nC esti Save id has a van EPIC
Answer #1

We know that,

Force,F = ma = q^2/(4*pi*Eo*r^2)

1*10^-3 kg * 414 m/s^2 = q^2/(4*pi*8.85*10^-12*0.02^2)

q = sqrt(0.414*4*pi*8.85*10^-12*0.02^2)

= 1.356*10^-7 C

= 135.6 nC

Option b is correct.